# On the Brun–Titchmarsh theorem

### v1

@@ 7.3. Estimates for exponential sums in $\mathcal{R}^{*}_{2}$
$\mathcal{R}^{*}_{2}(\cdots)$ $=\frac{\delta}{4\pi^{2}}\iint_{\mathbf{R}^{2}}\widetilde{F}(it_{1},it_{2})\mathrm{d}t_{1}\mathrm{d}t_{2}\int_{\frac{dN}{4MR^{\prime}\delta}}^{\frac{4dN}{MR^{\prime}\delta}}\mathrm{d}y\int_{U}^{V_{1}}y_{1}^{-it_{1}}\mathrm{d}y_{1}\int_{U}^{V_{2}}y_{2}^{-it_{2}}\mathrm{d}y_{2}$
×$\ \ \ \ \times\sum_{1\leqslant|h|\leqslant H\delta/d}\sum_{|r|\sim R^{\prime}}\mathrm{e}(-hry)r^{i(t_{1}+t_{2})-1}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}m_{1}\sim A_{1},~{}n_{1}\sim B_{1},~{}m_{2}\sim A_{2},~{}n_{2}\sim B_{2}\\ (m_{1}m_{2}n_{1}n_{2},q)=(m_{1}n_{1},m_{2}n_{2})=1\end{subarray}}\alpha^{\prime}_{d_{1}m_{1}}\alpha^{\prime\prime}_{d_{2}n_{1}}\beta^{\prime}_{d_{1}m_{2}}\beta^{\prime\prime}_{d_{2}n_{2}}$
@@
$\iint_{\mathbf{R}^{2}}|\widetilde{F}(it_{1},it_{2})|\mathrm{d}t_{1}\mathrm{d}t_{2}\ll(\eta N\log N)^{2}.$
@@
$(a,h,m,n,r,s)\leftarrow(q/\delta,k,m_{1},n_{1},m_{2},n_{2}),\ \ \ (H,M,N,R,S)\leftarrow(R^{\prime}H\delta/d,A_{1},B_{1},A_{2},B_{2}),$
$\alpha_{m,r}\leftarrow\alpha^{\prime}_{d_{1}m_{1}}\beta^{\prime}_{d_{1}m_{2}}m_{1}^{i(t_{2}-t_{1})}\mathrm{e}\Big{(}m_{1}m_{2}y_{1}y\delta-\frac{qy_{2}}{m_{1}m_{2}}\Big{)},$
$\beta_{h,n,s}\leftarrow\alpha^{\prime\prime}_{d_{2}n_{1}}\beta^{\prime\prime}_{d_{2}n_{2}}n_{2}^{i(t_{1}-t_{2})}\mathrm{e}(n_{1}n_{2}y_{2}y\delta)\mathop{\sum\sum}_{\begin{subarray}{c}hr=k\\ 1\leqslant|h|\leqslant H\delta/d,~{}|r|\sim R^{\prime}\end{subarray}}\mathrm{e}(-hry)r^{i(t_{1}+t_{2})-1},$
@@
$\mathcal{R}^{*}_{2}$ $\ll(MNq)^{\varepsilon}\frac{\eta^{4}N}{M}(RHM^{2})^{\frac{1}{2}}((RHB_{1}B_{2})^{\frac{1}{4}}+(B_{1}B_{2})^{\frac{1}{2}})$
×$\ \ \ \ \times\Bigg{[}q^{\vartheta}B_{1}B_{2}(A_{1}A_{2})^{\frac{1}{2}}\Big{(}1+\frac{qRH}{M^{2}}+\frac{A_{2}}{MB_{1}}\Big{)}^{\frac{1}{2}}\Big{(}1+\frac{qRH}{M^{2}}+\frac{RH}{B_{1}B_{2}}\Big{)}^{\frac{1}{2}}+B_{2}^{-2}A_{1}(RHB_{1}B_{2})^{\frac{1}{2}}\Bigg{]}^{\frac{1}{2}}$
$\ \ \ \ +\eta^{4}NM^{-1}(A_{1}A_{2}RHM^{2})^{\frac{1}{2}+\varepsilon}$
$\ll(MNq)^{\varepsilon}\eta^{\frac{5}{2}}MNq^{-\frac{1}{2}}\Big{(}\Big{(}\frac{M^{2}B_{1}B_{2}}{q}\Big{)}^{\frac{1}{4}}+(B_{1}B_{2})^{\frac{1}{2}}\Big{)}$
×$\ \ \ \ \times\Bigg{[}q^{\vartheta}B_{1}B_{2}(A_{1}A_{2})^{\frac{1}{2}}\Big{(}1+\frac{M^{2}}{B_{1}B_{2}q}\Big{)}^{\frac{1}{2}}+B_{2}^{-2}A_{1}\Big{(}\frac{M^{2}B_{1}B_{2}}{q}\Big{)}^{\frac{1}{2}}\Bigg{]}^{\frac{1}{2}}$
$\ \ \ \ +(MNq)^{\varepsilon}\eta^{\frac{5}{2}}MN(A_{1}A_{2}/q)^{\frac{1}{2}},$
Taking $B_{1}B_{2}=(1+q^{-1}M^{2})(MNq)^{4\varepsilon}$, we get
$\mathcal{R}^{*}_{2}$ $\ll(MNq)^{\varepsilon}\eta^{\frac{9}{2}}MNq^{-\frac{1}{2}}(1+Mq^{-\frac{1}{2}})[q^{\vartheta}B_{1}B_{2}(A_{1}A_{2})^{\frac{1}{2}}]^{\frac{1}{2}}$
$\ \ \ \ +(MNq)^{\varepsilon}\eta^{\frac{5}{2}}MN(A_{1}A_{2}/q)^{\frac{1}{2}}$
$\ll\eta^{\frac{9}{2}}M^{\frac{3}{2}}Nq^{\frac{\vartheta-1}{2}}(1+Mq^{-\frac{1}{2}})^{\frac{3}{2}}(MNq)^{3\varepsilon}+\eta^{\frac{9}{2}}MN(MNq)^{-\varepsilon},$
from which and (7.8), (7.7), (7.5), we then conclude the proof of Theorem 3.1 by noticing the arbitrariness of $\varepsilon.$
@@ 8.1. Brun–Titchmarsh theorem on average
Hooley [Ho72] was the first who was able to study upper bounds of $\pi(x;q,a)$ non-tivially through extra averaging over $q$. This was largely motivated by the study of greatest prime factors of shifted primes, as approximations to the twin prime conjecture. For an individual $q$, one may also consider the average over the residue class $a\,(\operatorname{mod}{q})$. See also [Mo74, Ho75, Mo79, Iw82, Fo85] for these developments.
8.2. A variant of Theorem 3.1
As one may see from the proof of Theorem 3.1, one needs to estimate the exponential sum
$\sum_{k\leqslant K}\delta_{k}\sum_{m\leqslant M}\sum_{n\leqslant N}\alpha_{m}\beta_{n}\mathrm{e}\Big{(}ak\frac{\overline{m}}{n}\Big{)}.$
Since the above $\alpha_{m},\beta_{n}$ both factorize into convolutions in the prerequisites of Theorem 3.1, it is in fact reasonable to bound the quintilinear form (7.1). The flexibilities in convolutions are quite crucial in the proof of Theorem 3.1, and this structure can be guaranteed by Lemma 2.1 as developed by Iwaniec.
For arbitrary coefficients $-β\boldsymbol{\alpha},\bbeta,$ Bettin and Chandee [BC18] obtained an estimate for the above trilinear form, as a successor of a bound due to Duke, Friedlander and Iwaniec [DFI97] with $K=1$.
Lemma 8.1.
[…]
×$\qquad\times((M+N)^{\frac{1}{4}}+(KMN)^{\frac{1}{40}}(KM+KN)^{\frac{1}{8}}).$
The arguments in Section 7, together with Lemma 8.1, yield the following alternative formulation of (3.2) for arbitrary coefficients.
Theorem 8.1.
[…]
where the implied constant depends only on $\varepsilon$.
We also recall that Friedlander and Iwaniec [FI92] studied the twisted second moment
$\mathcal{Z}:=\frac{1}{\varphi(q)}\sum_{\chi\,(\operatorname{mod}{q})}\Big{|}\sum_{m\leqslant M}c_{m}\chi(m)\Big{|}^{2}\Big{|}\sum_{l\leqslant L}\chi(l)\Big{|}^{2}$
with $|c_{m}|\ll m^{\varepsilon},$ which is intimately related to the counting function (3.1). We now compare [FI92], Theorems 3.1 and 8.1, with an expectation that an asymptotic formula holds for $\mathcal{Z}$ with $M$ as large as possible. Take $L=q^{1/2}$ for interest.
  • Friedlander and Iwaniec [FI92] succeeded for $M\leqslant q^{2/7-}$ and $(c_{m})$ is a convolution of two coefficients of level $q^{1/7-}.$
  • Theorem 3.1 succeeds for $M\leqslant q^{(3-2\vartheta)/4-}$ and $c_{m}$ is well-factorable.
  • Theorem 8.1 succeeds for $M\leqslant q^{2/3-}$ and $c_{m}$ is general.
Recall the quintilinear form in (7.1):
@@
$-β\displaystyle\mathcal{Q}_{a}(\boldsymbol{\alpha},\bbeta)$ $=\sum_{h\leqslant H}\mathop{\sum_{m\leqslant M}\sum_{n\leqslant N}\sum_{r\leqslant R}\sum_{s\leqslant S}}_{(mn,rs)=(ns,a)=1}\alpha_{m,r}\beta_{h,n,s}\mathrm{e}\Big{(}\frac{ah\overline{mn}}{rs}\Big{)},$
@@
which appears in the proofs of Theorem 3.1 and [FI92, Theorem 1]. In particular, due to the application of Kloostermania, the size of $a$ plays an important role in the settings of Theorem 3.1. Note that Lemma 8.1 follows from a careful application of amplification, so that only elementary means together with Weil’s bound for complete algebraic exponential sums are needed. A natural problem is to ask for an elementary bound for $-β\mathcal{Q}_{a}(\boldsymbol{\alpha},\bbeta)$ which is uniform in $a\neq 0$, and $-β\boldsymbol{\alpha},\bbeta$ could be well-factorable if necessary.
8.3. Twisted moments of Dirichlet $L$-functions
As one may see from the above arguments, it is important to study the twisted moment
$\frac{1}{\varphi(q)}\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}\Big{|}\sum_{m\leqslant M}c_{m}\chi(m)\Big{|}^{2}|L(\tfrac{1}{2}+it,\chi)|^{2k},$
and prove an averaged version of Lindelöf Hypothesis with $M$ as large as possible. We look forward to more advanced techniques to enter the picture. For $t=0,k=2$ and $M=q^{1/4}$ with $q$ prime, we refer to the work of Blomer, Humphries, Khan and Milinovich [BHKM20]. This is very cogent to make us believe that spectral reciprocity contributes as a very key input, so that the resultant bound is independent from the Ramanujan–Petersson conjecture. As mentioned before, it is highly desired to extend their bound to arbitrary moduli $q$.
@@ 8.4. Hooley’s Hypothesis R and impacts
are admissible in (1.3).
Hypothesis R exhibits squareroot cancellations within incomplete Kloosterman sums. It seems far beyond the current technique to prove Conjecture 8.1 for general $q$, even in the case $|I|\approx q^{0.99}.$ On the other hand, Conjecture 8.1 corresponds to the exponent pair hypothesis in the theory of arithmetic exponent pairs, for which we hope $(\kappa,\lambda)=(0,\frac{1}{2})$ is admissible for (4.2). This illustrates that all of our previous arguments should never beat (8.1) unconditionally even imposing factorization conditions on $q$.
However, as $\varpi\rightarrow 9/20-$ for instance, Theorem 1.5 shows that $C(\varpi)=5/(5-6\varpi)$ is admissible, as long as $q$ is squarefree and smooth. This is obviously sharper than the conjectural bound 8.1 which takes $C(\varpi)=6/(5-6\varpi).$ As one may see from previous sections, Theorem 1.5 is proved by using orthogonality of multiplicative characters, so that Lemma 4.6 on character sums has been applied to finish the proof. This phenomenon indicates that for $\varpi<1/2-$ it is clever to use multiplicative characters, and for $\varpi>1/2+$ additive ones might be more powerful instead. The exact transition range is very mysterious to the authors.

### v3

@@ 7.3. Estimates for exponential sums in $\mathcal{R}^{*}_{2}$
$\mathcal{R}^{*}_{2}(\cdots)$ $=\frac{\delta}{4\pi^{2}}\iint_{\mathbf{R}^{2}}\widetilde{F}(it_{1},it_{2})\mathrm{d}t_{1}\mathrm{d}t_{2}\int_{\frac{L}{8A_{2}B_{2}R^{\prime}\delta}}^{\frac{2L}{A_{2}B_{2}R^{\prime}\delta}}\mathrm{d}y\int_{U}^{V_{1}}y_{1}^{-it_{1}}\mathrm{d}y_{1}\int_{U}^{V_{2}}y_{2}^{-it_{2}}\mathrm{d}y_{2}$
×$\ \ \ \ \times\sum_{1\leqslant|h|\leqslant H}\sum_{|r|\sim R^{\prime}}\mathrm{e}(-hry)r^{i(t_{1}+t_{2})-1}$
×$\ \ \ \ \times\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}m_{1}\sim A_{1},~{}n_{1}\sim B_{1},~{}m_{2}\sim A_{2},~{}n_{2}\sim B_{2}\\ (m_{1}m_{2}n_{1}n_{2},q)=(m_{1}n_{1},m_{2}n_{2})=(m_{1},d_{2})=(m_{2},d_{2}^{\prime})=1\end{subarray}}\alpha^{\prime}_{d_{1}m_{1}}\alpha^{\prime\prime}_{d_{2}n_{1}}\beta^{\prime}_{d_{1}^{\prime}m_{2}}\beta^{\prime\prime}_{d_{2}^{\prime}n_{2}}$
@@
$\iint_{\mathbf{R}^{2}}|\widetilde{F}(it_{1},it_{2})|\mathrm{d}t_{1}\mathrm{d}t_{2}\ll(\eta L\log L)^{2}.$
@@
$(a,h,m,n,r,s)\leftarrow(q/\delta,k,m_{1},n_{1},m_{2},n_{2}),\ \ \ (H,M,N,R,S)\leftarrow(R^{\prime}H,A_{1},B_{1},A_{2},B_{2}),$
$\alpha_{m,r}\leftarrow\alpha^{\prime}_{d_{1}m_{1}}\beta^{\prime}_{d_{1}^{\prime}m_{2}}m_{1}^{i(t_{2}-t_{1})}\mathrm{e}\Big{(}m_{1}m_{2}y_{1}y\delta-\frac{qy_{2}}{m_{1}m_{2}}\Big{)},$
$\beta_{h,n,s}\leftarrow\alpha^{\prime\prime}_{d_{2}n_{1}}\beta^{\prime\prime}_{d_{2}^{\prime}n_{2}}n_{2}^{i(t_{1}-t_{2})}\mathrm{e}(n_{1}n_{2}y_{2}y\delta)\mathop{\sum\sum}_{\begin{subarray}{c}hr=k\\ 1\leqslant|h|\leqslant H,~{}|r|\sim R^{\prime}\end{subarray}}\mathrm{e}(-hry)r^{i(t_{1}+t_{2})-1},$
@@
$\mathcal{R}^{*}_{2}$ $\ll(qNL)^{\varepsilon}\eta^{4}L(RHN)^{\frac{1}{2}}((RHN_{1}N_{2})^{\frac{1}{4}}+(N_{1}N_{2})^{\frac{1}{2}})$
×$\ \ \ \ \times\Bigg{[}q^{\vartheta}(N_{1}N_{2})^{\frac{1}{2}}\Big{(}1+\frac{qRH}{N^{2}}\Big{)}^{\frac{1}{2}}\Big{(}1+\frac{qRH}{N^{2}}+\frac{RH}{N_{1}N_{2}}\Big{)}^{\frac{1}{2}}\Bigg{]}^{\frac{1}{2}}+\eta^{4}NL(RH/N_{1}N_{2})^{\frac{1}{2}+\varepsilon}$
$\ll\eta^{\frac{9}{2}}N^{\frac{3}{2}}Lq^{\frac{\vartheta-1}{2}}(N_{1}N_{2})^{\frac{1}{2}}(N^{2}q^{-1}+N_{1}N_{2})^{\frac{1}{4}}\Big{(}1+\frac{N^{2}}{qN_{1}N_{2}}\Big{)}^{\frac{1}{4}}(qNL)^{\varepsilon}+\eta^{\frac{9}{2}}N^{2}L(qN_{1}N_{2})^{-\frac{1}{2}}(qNL)^{\varepsilon},$
from which and (7.8), (7.7), we then conclude the proof of Theorem 2.2.
Moreover, in the settings of Theorem 2.1, the coefficients $-β\boldsymbol{\alpha},\bbeta$ are arbitrary, and we appeal to Lemma 2.3 instead to bound the sums with Kloosterman fractions in (7.9). To this end, one may argue as above to group the variables, and the details are omitted here.
@@ 8. Comments and remarks
8.1. Twisted fourth moment of Dirichlet $L$-functions
(8.1) $\frac{1}{\varphi(q)}\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|L(\tfrac{1}{2}+it,\chi)|^{4}\ll q^{\varepsilon}(1+|t|)^{O(1)}$
with $\delta>0$ as large as possible.
By extending Motohashi’s identity to link fourth moment of Dirichlet $L$-functions and cubic moment of $\mathrm{GL}_{2}$ $L$-functions, Blomer, Humphries, Khan and Milinovich [BHKM20] prove that (8.2) is valid as long as $q$ is prime, $t=0$, $\delta\in[0,1/4]$ and $\boldsymbol{\alpha}$ is supported on squarefree numbers. Their bound can, together with a suitable amplification, lead to a new proof of Burgess’s bound $L(\frac{1}{2},\chi)\ll q^{3/16+\varepsilon}$ for prime $q$. In applications to Brun–Titchmarsh theorem, it is highly desirable to extend (8.2) for a larger admissible $\delta.$ Also, it is desirable to extend [BHKM20] to all moduli $q$, which then readily implies that the choice (1.6) is admissible in (1.4) for all such $q$.
Recall that we have utilized (8.1) in the proof of Theorem 1.5. If one was allowed to use (8.2) with smooth and squarefree $q$, it would be possible to obtain the desirable constant $2$ in Theorem 1.5 in a larger range of $\varpi$.
@@ 8.2. Brun–Titchmarsh theorem on average
Hooley [Ho72] was the first who was able to study upper bounds of $\pi(x;q,a)$ non-tivially through extra averaging over $q$. This was largely motivated by the study of greatest prime factors of shifted primes, as approximations to the twin prime conjecture. For an individual $q$, one may also consider the average over the residue class $a\,(\operatorname{mod}{q})$. See also [Mo74, Ho75, Mo79, Iw82, Fo84b, Fo85, BH96, FR22] for these developments.
8.3. Extending Theorem 1.2 to general moduli
In his ETH thesis, Löffel [Lo16] generalized Lemma 3.3, as well as other works of Fouvry, Kowalski and Michel [FKM14, FKM15], to squarefree moduli. This also allows us to relax $q$ in Theorem 1.2 to be squarefree. It is also possible to generalize Theorem 1.2 to arbitrary moduli along with the same line. As one may see from the proof of Lemma 3.3, it is necessary to evaluate averages of products of four Kloosterman sums to general moduli.
In another direction, Lin and Michel [LM24] considered a variant of Lemma 3.3 with modulus $q$, which is exactly a product of two distinct primes. When the two prime factors are of suitable sizes, they are able to capture stronger non-correlations between algebraic trace functions and Hecke eigenvalues (holomorphic or Maass). By extending their arguments to the situation of Eisenstein spectrum, it is also possible to study the Brun–Titchmarsh theorem with factorable moduli, which can be compared with the direct application of the method of arithmetic exponent pairs.
8.4. Twisted second moment of character sums
Recall that Theorems 2.1, 2.2 and Lemmas 2.5, 2.6 are intimately related to the twisted second moment of character sums
$\mathcal{Z}:=\frac{1}{\varphi(q)}\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}\Big{|}\sum_{m\leqslant M}c_{m}\chi(m)\Big{|}^{2}\Big{|}\sum_{l\leqslant L}\chi(l)\Big{|}^{2}$
with $|c_{m}|\ll m^{\varepsilon}.$ We expect that $\mathcal{Z}\ll MLq^{\varepsilon}$ holds with $M$ as large as possible, building on the squareroot cancellation philosophy.
  • Friedlander and Iwaniec [FI92] succeeded for $M\leqslant q^{1/2+1/22-}$ and $(c_{m})$ is a convolution of two coefficients of level $q^{3/11-}.$
  • Duke, Friedlander and Iwaniec [DFI97a] succeeded for $M\leqslant q^{1/2+1/190-}$ and $c_{m}$ is general.
  • Lemma 2.5 succeeds for $M\leqslant q^{1/2+1/66-}$ and $c_{m}$ is general.
  • Lemma 2.6 succeeds for $M\leqslant q^{(5-2\vartheta)/8-}$ and $c_{m}$ is well-factorable.
Recall that Theorem 2.1 and Lemma 2.5 are proven with the aid of Lemma 2.3 on trilinear forms wiht Kloosterman fractions. However, Theorem 2.2 and Lemma 2.6 rely on the control of the quintilinear form in (7.1):
@@ 8.4. Twisted second moment of character sums
$-β\displaystyle\mathcal{Q}_{a}(\boldsymbol{\alpha},\bbeta)$ $=\sum_{h\leqslant H}\mathop{\sum_{m\leqslant M}\sum_{n\leqslant N}\sum_{r\leqslant R}\sum_{s\leqslant S}}_{(mn,rs)=(ns,a)=1}\alpha_{m,r}\beta_{h,n,s}\mathrm{e}\Big{(}\frac{ah\overline{mn}}{rs}\Big{)}.$
@@
Due to the application of Kloostermania, the size of $a$ plays an important role in the settings of Theorem 2.2. Note that Lemma 2.3 follows from a careful application of amplification, so that only elementary means together with Weil’s bound for complete algebraic exponential sums are needed. A natural problem is to ask for an elementary bound for $-β\mathcal{Q}_{a}(\boldsymbol{\alpha},\bbeta)$ which is uniform in $a\neq 0$, and $-β\boldsymbol{\alpha},\bbeta$ are of convolution structures if necessary.
@@ 8.5. Hooley’s Hypothesis R and impacts
are admissible in (1.3). Hypothesis R exhibits squareroot cancellations within incomplete Kloosterman sums. It seems far beyond the current technique to prove Conjecture 8.1 for general $q$, even in the case $|I|\approx q^{0.99}.$ On the other hand, Conjecture 8.1 corresponds to the exponent pair hypothesis in the theory of arithmetic exponent pairs, for which we hope $(\kappa,\lambda)=(0,\frac{1}{2})$ is admissible for (3.2). This illustrates that all of our previous arguments starting from additive characters should never beat (8.3) unconditionally even imposing factorization conditions on $q$.
However, as $\varpi\rightarrow 9/20-$ for instance, Theorem 1.5 shows that $C(\varpi)=5/(5-6\varpi)$ is admissible, as long as $q$ is squarefree and smooth. This is obviously sharper than the conjectural bound (8.3) which takes $C(\varpi)=6/(5-6\varpi).$ As one may see from previous sections, Theorem 1.5 is proved by using orthogonality of multiplicative characters, so that Lemma 3.6 on character sums has been applied to finish the proof. This phenomenon indicates that for $\varpi<1/2-$ it is clever to use multiplicative characters, and for $\varpi>1/2+$ additive ones might be more powerful instead. The exact transition range is very mysterious to the authors.